Worth 5 points (Science: day of the week) Zellers congruence is an algorithm developed by Christian Zeller to calculate thein python
Answer #1

Program to implement Zeller's Congruence.

Python Code:

# Program to implement Zeller's Congruence.
# Written in Python 3

#Take input from the user
y = int(input("Enter year: (e.g., 2008): "))

m = int(input("Enter month: 1-12: "))
if m==1:
    m=13
    y=y-1
elif m==2:
    m=14
    y=y-1
# extra processing for Jan and Feb, we also subtract one from year

q = int(input("Enter the day of month: 1-31: "))

j = y//100 
k = y%100

h = (q +(26*(m+1))//10 +k +k//4 +j//4 +(5*j)) % 7
#calculate the day of week by the Zeller's Congruence

week =["Saturday","Sunday","Monday","Tuesday","Wednesday","Thursday","Friday"] #listofweeks
print("Day of week is: ", week[h] )
#print the result


Screenshot of Output:

Shell Enter year: (e.g., 2008): 2013 Enter month: 1-12: 1 Enter the day of month: 1-31: 25 Day of week is: Friday

Sample inputs are mentioned in question.